OpenStudy (anonymous):

A rabbit and a turtle were to start a race at the same place. The rabbit was asleep when the turtle started. He slept for seven hours before he started the race. The rabbit's average rate when he was awake was 5 km/hr. The turtle never stopped for a break, but his pace was only 1km/hr. How long will the rabbit run before it catches the turtle?

OpenStudy (anonymous):

First find how much distance the turtle would have covered in 7 hours.

OpenStudy (anonymous):

How much distance?

OpenStudy (anonymous):

turttle travelled distance = 1(7+t), where t is time from when rabit wake up when rabit reaches turtle, their distances will be equal: 1*(7+t) = 5*t solve this for t